# AVadCLIP: Audio-Visual Collaboration for Robust Video Anomaly Detection

### v1

@@ 4.4 Ablation Studies
Table 4: Effectiveness of designed modules on XD-Violence.
@@
Table 5: Effectiveness of audio-visual fusion on XD-Violence.
@@
Table 6: Effectiveness of UKD on XD-Violence.
The effect of audio-visual adaptive fusion
@@
From Table 4, it can be observed that the introduction of audio-visual fusion improves detection performance. Furthermore, Table 5 presents the impact of different audio-visual fusion strategies on anomaly detection performance. First, the cross attention fusion performs poorly in the WSVAD task, indicating that although it can capture the relationships between modalities, its complex parameterized design may negatively impact the generalization ability of CLIP model in downstream WSVAD tasks. Next, the simple element-wise addition strategy achieves an AP of 83.02% and an AVG of 27.66%. Then, the concatenation with linear projection approach improves the AP to 83.36% and the AVG to 28.88%, indicating that enhancing feature representation through linear transformation facilitates more effective cross-modal information capture. Finally, our proposed adaptive fusion strategy achieves the best AP of 86.04%, outperforming the other three methods on the whole. This demonstrates that our adaptive fusion strategy, as a lightweight and effective fusion strategy, can more exploit complementary information between audio and visual modalities.
@@
Figure 4: Coarse- and Fine-grained WSVAD visualization results of AVadCLIP and Baseline on XD-Violence.
The effect of audio-visual prompt and $\mathcal{L}_{FOCAL}$
@@
As presented in Table 4, the baseline model achieves an AP of only 79.85%. Integrating the audio-visual prompt on top of the adaptive fusion mechanism significantly enhances performance, increasing the AP to 86.18%. This improvement underscores the effectiveness of the audio-visual prompt in capturing critical multimodal patterns, thereby facilitating more precise anomaly recognition. Furthermore, incorporating focal loss into the model contributes to refining anomaly boundary detection, leading to more stable performance in fine-grained anomaly localization. In summary, the audio-visual prompt primarily enhances coarse-grained anomaly detection, and focal loss further refines boundary precision, enabling the model to achieve optimal performance across both AP and AVG metrics.
The effect of uncertainty-driven distillation
@@
As shown in Table 6, the proposed UKD mechanism significantly enhances anomaly detection performance in both visual-only and audio-only models. Specifically, in the visual-only setting, UKD achieves a 0.9% improvement in AP and a 4.5% increase in AVG, attaining performance levels comparable to the teacher model trained with audio-visual inputs. Similarly, the audio-only model also benefits from UKD, exhibiting consistent performance gains. These results highlight the effectiveness of UKD in leveraging data uncertainty to enhance the robustness of unimodal representations during the distillation process, making it particularly well-suited for real-world applications where modality incompleteness is prevalent.
4.5 Qualitative Results
@@
In Figure 4, we present the qualitative visualizations of AVadCLIP and the baseline model for both coarse-grained and fine-grained WSVAD. The blue curves denote the anomaly predictions by AVadCLIP, whereas the yellow curves represent those by the baseline model. As illustrated, compared to the baseline model, AVadCLIP significantly reduces anomaly confidence in normal video segments, thereby enhancing its ability to distinguish between abnormal and normal regions more accurately. The fine-grained map below also indicates that AVadCLIP can predict categories with greater precision. Notably, the observed performance improvement supports our hypothesis that audio information is more advantageous in visual occlusion (shooting) or acoustic dominant scenes (explosion), and can effectively eliminate ambiguity in visually similar patterns in anomaly detection scenes, thereby ensuring more robust detection performance.
5 Conclusion

### v2

@@ IV-D Ablation Studies
TABLE VII: Effectiveness of audio-visual fusion on XD-Violence.
@@
TABLE VIII: Effectiveness of UKD on XD-Violence.
@@
TABLE IX: Effectiveness of UKD on CCTV-Fightssub.
Method AP(%) AVG(%)
Visual Model w/o UKD 67.89 10.65
Visual Model w UKD 73.36 11.02
Audio-Visual Model 73.38 12.17
IV-D1 The effect of audio-visual adaptive fusion
@@
From Table VI, it can be observed that the introduction of audio-visual fusion improves detection performance. Furthermore, Table VII presents the impact of different audio-visual fusion strategies on anomaly detection performance. First, the cross attention fusion performs poorly in the WSVAD task, indicating that although it can capture the relationships between modalities, its complex parameterized design may negatively impact the generalization ability of CLIP model in downstream WSVAD tasks. Next, the simple element-wise addition strategy achieves an AP of 83.02% and an AVG of 27.66%. Then, the concatenation with linear projection approach improves the AP to 83.36% and the AVG to 28.88%, indicating that enhancing feature representation through linear transformation facilitates more effective cross-modal information capture. Finally, our proposed adaptive fusion strategy achieves the best AP of 86.04%, outperforming the other three methods on the whole. This demonstrates that our adaptive fusion strategy, as a lightweight and effective fusion strategy, can more exploit complementary information between audio and visual modalities.
@@
Figure 4: Coarse-grained and Fine-grained WSVAD visualization results of AVadCLIP and the baseline model on XD-Violence.
Figure 5: Coarse-grained WSVAD visualization results of AVadCLIP, AVadCLIP, and the baseline model on XD-Violence.
Figure 6: Coarse-grained WSVAD visualization results of AVadCLIP, AVadCLIP, and the baseline model on CCTV-Fightssub.
IV-D2 The effect of audio-visual prompt and $\mathcal{L}_{FOCAL}$
@@
As presented in Table VI, the baseline model achieves an AP of only 79.85%. Integrating the audio-visual prompt on top of the adaptive fusion mechanism significantly enhances performance, increasing the AP to 86.18%. This improvement underscores the effectiveness of the audio-visual prompt in capturing critical multimodal patterns, thereby facilitating more precise anomaly recognition. Furthermore, incorporating focal loss into the model contributes to refining anomaly boundary detection, leading to more stable performance in fine-grained anomaly localization. In summary, the audio-visual prompt primarily enhances coarse-grained anomaly detection, and focal loss further refines boundary precision, enabling the model to achieve optimal performance across both AP and AVG metrics.
IV-D3 The effect of uncertainty-driven distillation
@@
As shown in Table VIII, the proposed UKD mechanism significantly enhances anomaly detection performance in both visual-only and audio-only models. Specifically, in the visual-only setting, UKD achieves a 0.9% improvement in AP and a 4.5% increase in AVG, attaining performance levels comparable to the teacher model trained with audio-visual inputs. Similarly, the audio-only model also benefits from UKD, exhibiting consistent performance gains. These results highlight the effectiveness of UKD in leveraging data uncertainty to enhance the robustness of unimodal representations during the distillation process, making it particularly well-suited for real-world applications where modality incompleteness is prevalent. In addition, we present a comparison of the effectiveness of the UKD module on CCTV-Fightssub in Table IX. The results show that the proposed UKD mechanism significantly improves the anomaly detection performance of the unimodal model. Notably, in the visual-only scenario, adding UKD improves AP by 5.5%, achieving performance comparable to that of the audio-visual model. This finding further demonstrates the effectiveness of UKD in enhancing the robustness of unimodal representations through data uncertainty.
IV-E Qualitative Results
@@
In Figure 4, we present the qualitative visualizations of AVadCLIP and the baseline model for both coarse-grained and fine-grained WSVAD. The blue curves denote the anomaly predictions by AVadCLIP, whereas the yellow curves represent those by the baseline model (RGB-only w/o UKD). As illustrated, compared to the baseline model, AVadCLIP significantly reduces anomaly confidence in normal video segments, thereby enhancing its ability to distinguish between abnormal and normal regions more accurately. The fine-grained map below also indicates that AVadCLIP can predict categories with greater precision. Notably, the observed performance improvement supports our hypothesis that audio information is more advantageous in visual occlusion (shooting) or acoustic dominant scenes (explosion), and can effectively eliminate ambiguity in visually similar patterns in anomaly detection scenes, thereby ensuring more robust detection performance.
In addition, we compare the coarse-grained visualization results of the baseline model (RGB-only w/o UKD), the student model AVadCLIP, and the teacher model AVadCLIP on XD-Violence, as shown in Figure 5. Experimental results show that AVadCLIP significantly outperforms the other two counterparts. By using this model as a teacher model to guide the unimodal student model, it effectively mitigates anomaly confidence biases, steering them towards more accurate detection results. To a certain extent, this demonstrates the robustness of our proposed method.
In order to demonstrate the superiority of our proposed method more comprehensively and intuitively, Figure 6 shows the coarse-grained visualization results on the CCTV-Fightssub dataset (since this dataset only includes the “Fighting” category, fine-grained visualizations are not provided). It can be seen that our method achieves significantly higher anomaly confidence scores in abnormal regions and notably lower scores in normal regions compared to the baseline model. This demonstrates that the integration of audio and video information can still yield substantial performance improvements in complex scenes. Besides, as can be seen from the last two rows, the unimodal model distilled with the UKD mechanism shows significantly fewer false positives compared to the baseline, demonstrating that the UKD mechanism effectively transfers audio-visual multi-modal knowledge into the unimodal model.
V Conclusion